OpenStudy (anonymous):

\[\sqrt{5}(10 - 4 \sqrt{2})\]

OpenStudy (anonymous):

Distributive Law!

OpenStudy (anonymous):

distribute the sqrt(5)

OpenStudy (anonymous):

\[10\sqrt{5}-4\sqrt{10}\]

jhonyy9 (jhonyy9):

do you like factorizing out the 2sqrt5 ? do you know it how ?

jhonyy9 (jhonyy9):

what you need calcule from this again ?

OpenStudy (anonymous):

I got \[5\sqrt{2} - 4\sqrt{10}\]